OpenStudy (anonymous):

Hello, I need help with a problem in surface area of prisms and cylinders. Can anyone help.

OpenStudy (beasturias):

hello

OpenStudy (anonymous):

i can try

OpenStudy (beasturias):

i can try to

OpenStudy (anonymous):

okay, thanks. Here is the problem. find the lateral area for the given prism.

OpenStudy (beasturias):

what prism

OpenStudy (anonymous):

This is the one. Can you help please!

OpenStudy (anonymous):

do u have choices?

OpenStudy (anonymous):

anyways you find the area of one side and multiply it by 2 bcuz there are 2 of the same sides, and then do that to all of the sides and add them together 6*26*2=312 15*26*2=780 6*15*2=180 add them all up and you get 1,272 and that is ur anwser

OpenStudy (anonymous):

@SummerMoriarty3 did this help?

OpenStudy (anonymous):

Yes, thank you so much fro helping. Could you help me with another one please?

OpenStudy (anonymous):

Find the area of the given cylinder. Use 3.14 for pi and round to the nearest whole number.

OpenStudy (anonymous):

sure np

OpenStudy (anonymous):

433.54

OpenStudy (anonymous):

The answer choices are A. 180 ft^2 B. 720 ft^2 C. 433 ft^2 D. 135 ft^2

OpenStudy (anonymous):

So, it would be C. ?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

do u have anymore?

OpenStudy (anonymous):

Thanks so much. 2 more okay. Is that okay with you?

OpenStudy (anonymous):

yea thats fine :)

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

Find the lateral area for the given cylinder. use 3.14 and round to the nearest whole number.

OpenStudy (anonymous):

The answer choices are A. 144 yd^2 B. 288 yd^2 C. 2.712 yd^2 D. 904 yd^2

OpenStudy (anonymous):

umm i put it in google and i got 1130.97 i dont think thats right

OpenStudy (anonymous):

that is not one of my answer choices. But we have to move on. I have to leave in a couple of minutes.

OpenStudy (anonymous):

if thats okay with you?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Find the surface area for the following net.

OpenStudy (anonymous):

Answer choices are A. 184 cm^2 B. 200 cm^2 C. 120 cm^2 D. 174 cm^2

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

Thank you so much. I will tell you if it was right.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

number 2 was A instead of C. I got an 80%. Thanks so much for all your help. Bye

OpenStudy (anonymous):

Merry Christmas and a Happy New Year.

OpenStudy (anonymous):

You too bye!